Dividend from shares of Co. XYZ paid on 31.12.2018 was 1,000 PLN per share. It is expected that for the first three years (2019-2021) dividend will increase with a rate of 6% yearly and then in the following years it will increase with a rate of 4% yearly. The expected market rate is fixed in the whole investment period and equals 10%. Assuming that investor does not plan to sell shares of Co. XYZ what should be the market price according to investors prediction? Should investor purchase this share if the actual market price is 23,000 PLN per share?
Answer #1

Please provide rating..

we have to use 2 stage dividend discount model to solve this problem
as per DDM price of share is present value of future cash flow
i ii iii=i+ii iv v=iii*iv
Year Dividend Terminal value (note-1) Total cash flow PVIF @ 10% present value
1 =1000*106% 1,060.00           1,060     0.9091              964
2 =1000*106%*106% 1,123.60           1,124     0.8264              929
3 =1000*106%*106% 1,191.02          20,644.28        21,835     0.7513        16,405
       18,297
As per DDM we are getting price of share should be 18297 but actual market price is 23000
Therefore investor should not purchase share.
Note -1
Computation of terminal/horizon value
Price of share at the end of year 3 = D4/(required rate - growth rate)
=1191.02*104%/(10%-4%)
20644.28